[Hallc_running] [New Logentry] RC report Nov-27

berdnik at jlab.org berdnik at jlab.org
Sat Nov 27 09:55:02 EST 2021

Logentry Text:

9-14am Nov27: we are continuing to take data at the Q2=6.0, W=2.40, TH_shms=11.12deg (center) setting. We have 7543 pi- events at this point.
Day shift and swing shift yesterday smooth data taking.
At the end of swing we lost SHMS Q3 due communication error. Steve reset interlock remotely. The issue cost us ~45min of downtime.  https://logbooks.jlab.org/entry/3951284 and https://logbooks.jlab.org/entry/3951285
Start of the owl, we switched to dummy target and took all 3 hours of data for central setting. After dummy runs returned to physics production on LD2. 
The beam trip rate stays relatively high for the rest of owl and start of the day shift today.  


Run plan:

-- 11.12deg: 57hrs from Thursday at 22:30 is Sunday at 07:30, plus 3hrs of Dummy goes to Sunday at 10:30.
-- 13.12deg: 57hrs from Sunday at 10:30 is Tuesday at 19:30, plus 3hrs of Dummy goes to Tuesday at 22:30.

Pending issues:


RC meeting:

Let's have a meeting today at 10:00 AM EST.
Connection information: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3951497
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211127/05bb78bf/attachment.html>

More information about the Hallc_running mailing list